# Spectral-Operator Calculus (Part A): Trace-Form Evaluators and Spectral Growth Taxonomy

### v1

@@ B.3. Tensor-product bounds in the polynomial class
We now record the tensor-product estimates used in the polynomial part of Theorem~3.5. Let $D$ and $E$ be self-adjoint operators with nonnegative, discrete spectra and counting functions $N_{D},N_{E}$.
@@
and both counting functions are eventually nondecreasing, as in Convention~2.2.
Under these hypotheses $D$ and $E$ belong to the polynomial growth class $\mathcal{G}_{\mathrm{poly}}$ of Definition~3.4.
The eigenvalues of the tensor product $D\otimes E$ are the pairwise products $\lambda_{m}\mu_{n}$, where $\{\lambda_{m}\}$ and $\{\mu_{n}\}$ are the eigenvalues of $D$ and $E$, listed with multiplicity. A standard argument shows that the counting function of $D\otimes E$ can be written as a Stieltjes convolution:
$N_{D\otimes E}(\Lambda)=\sum_{m\geq 1}N_{E}\!\left(\frac{\Lambda}{\lambda_{m}}\right)=\int_{(0,\infty)}N_{E}\!\left(\frac{\Lambda}{t}\right)dN_{D}(t),\qquad\Lambda>0.$ (B.2)
@@
When $N_{D}(t)\sim C_{D}t^{d_{D}}$ with $C_{D}>0$, the above bound and its converse can be seen as a special case of Karamata’s Tauberian theorem for monotone functions; see Bingham-Goldie-Teugels [5, Theorem 1.7.1]. Here we only require the one-sided $O$-bound.
@@
Applying Corollary~B.8 under these hypotheses gives
@@
for every $\varepsilon>0$, so $D\otimes E\in\mathcal{G}_{\mathrm{poly}}$. This is precisely the tensor-product stability statement in part (3) of Theorem~3.5.
@@ C.1. Dropping dominated continuity
It is standard that such an $L$ exists for bounded real sequences.
@@ C.3. Summary
Thus the dominated continuity and projector-locality assumptions in Section 3.1 are not redundant: both are needed, in addition to unitary invariance, extensivity, and normalization, to obtain the rigidity result of Theorem~3.1.
End of Spectral-Operator Calculus (Part A)

### v2

@@ B.3. Tensor-product bounds in the polynomial class
We now record the tensor-product estimates used in the polynomial part of Theorem 3.5. Let $D$ and $E$ be self-adjoint operators with nonnegative, discrete spectra and counting functions $N_{D},N_{E}$.
@@
and both counting functions are eventually nondecreasing, as in Convention 2.2.
Under these hypotheses $D$ and $E$ belong to the polynomial growth class $\mathcal{G}_{\mathrm{poly}}$ of Definition 3.4.
The eigenvalues of the tensor product $D\otimes E$ are the pairwise products $a_{m}b_{n}$, where $\{a_{m}\}_{m\geq 1}$ and $\{b_{n}\}_{n\geq 1}$ are the eigenvalues of $D$ and $E$, listed with multiplicity. A standard argument shows that the counting function of $D\otimes E$ can be written as a Stieltjes convolution:
$N_{D\otimes E}(\Lambda)=\sum_{m\geq 1}N_{E}\!\left(\frac{\Lambda}{a_{m}}\right)=\int_{(0,\infty)}N_{E}\!\left(\frac{\Lambda}{t}\right)dN_{D}(t),\qquad\Lambda>0.$ (B.2)
@@
When $N_{D}(t)\sim C_{D}t^{d_{D}}$ with $C_{D}>0$, the above bound and its converse can be seen as a special case of Karamata’s Tauberian theorem for monotone functions; see Bingham-Goldie-Teugels [5]*Theorem 1.7.1. Here we only require the one-sided $O$-bound.
@@
Applying Corollary B.8 under these hypotheses gives
@@
for every $\varepsilon>0$, so $D\otimes E\in\mathcal{G}_{\mathrm{poly}}$. This is precisely the tensor-product stability statement in part (3) of Theorem 3.5.
@@ C.1. Dropping dominated continuity
It is standard that such an $L$ exists for bounded real sequences [10].
@@ C.3. Summary
Thus the dominated continuity and projector-locality assumptions in Section 3.1 are not redundant: both are needed, in addition to unitary invariance, extensivity, and normalization, to obtain the rigidity result of Theorem 3.1.